# From Roots to Rewards: Dynamic Tree Reasoning with RL

### v2

@@ Abstract
Modern language models address complex questions through chain-of-thought (CoT) reasoning Wei et al. (2023) and retrieval augmentation Lewis et al. (2021), yet struggle with error propagation and knowledge integration. Tree-structured reasoning methods, particularly the Probabilistic Tree-of-Thought (ProbTree)Cao et al. (2023) framework, mitigate these issues by decomposing questions into hierarchical structures and selecting answers through confidence-weighted aggregation of parametric and retrieved knowledge Yao et al. (2023). However, ProbTree’s static implementation introduces two key limitations: (1) the reasoning tree is fixed during the initial construction phase, preventing dynamic adaptation to intermediate results, and (2) each node requires exhaustive evaluation of all possible solution strategies, creating computational inefficiency.
We present a dynamic reinforcement learning Sutton and Barto (2018) framework that transforms tree-based reasoning into an adaptive process. Our approach incrementally constructs the reasoning tree based on real-time confidence estimates, while learning optimal policies for action selection (decomposition, retrieval, or aggregation). This maintains ProbTree’s probabilistic rigor while improving both solution quality and computational efficiency through selective expansion and focused resource allocation.
@@
From Roots to Rewards: Dynamic Tree Reasoning with RL
Ahmed Bahloul  and Simon Malberg School of Computation, Information and Technology Technical University of Munich, Germany {ahmed.bahloul, simon.malberg}@tum.de
1 Introduction
Large language models (LLMs) augmented with tree-structured reasoning, such as ProbTree Cao et al. (2023), have shown promise in answering knowledge-intensive complex questions by decomposing queries into sub-questions and probabilistically combining parametric and retrieved knowledge Petroni et al. (2019).
The ProbTree framework tackles complex question answering by constructing a tree-structured reasoning process. At the root of the tree, the input question is decomposed into smaller sub-questions, forming hierarchical nodes in the reasoning tree. At each node, the framework employs three possible solution strategies: (1) Closed-Book (CB), which leverages parametric knowledge from the language model alone; (2) Open-Book (OB), relying on external retrieval to augment the language model’s outputs; and (3) Child-aggregation (CA), where answers are recursively constructed by combining the outputs of child nodes in the tree. Each node computes a probabilistic confidence score for the outputs of these strategies, and the highest-confidence answer is selected. This selection process ensures that responses are grounded in both parametric (language model) and non-parametric (retrieved) knowledge, making ProbTree well-suited for handling knowledge-intensive queries. However, this static pipeline—where the full reasoning tree and strategy choices are decided in a single step—limits flexibility and introduces computational inefficiencies. Moreover, its brute-force strategy of evaluating all possible solutions per node (Closed-Book, Open-Book, and Child-aggregation) introduces computational inefficiency, especially for deep or wide trees. [^1] Code available at: https://github.com/ahmedehabb/From-Roots-to-Rewards-Dynamic-Tree-Reasoning-with-RL
1.1 Motivation
@@ 1.3 Contributions
  • Empirical validation showing competitive performance with reduced computational cost on standard benchmarks (HotpotQA Yang et al. (2018), MusiQue Trivedi et al. (2022), 2WikiMultihopQA Ho et al. (2020))
@@
2 Reinforcement Learning Framework for Dynamic Tree-of-Thought Reasoning
To overcome the limitations of static and brute-force tree generation in the original ProbTree framework Cao et al. (2023), we introduce a reinforcement learning (RL) based agent that dynamically generates reasoning trees and selects actions at each step to answer knowledge-intensive questions efficiently and accurately.
2.1 Problem Formulation
We formulate the decision process of tree construction and answer selection as a Markov Decision Process (MDP) Sutton and Barto (2018), where:
  • State ($s$): Encodes information about the current node in the tree, including semantic features, structural position, confidence scores, and answer embeddings Reimers and Gurevych (2019).
@@ 2.1 Problem Formulation
  • Reward ($r$): A combination of semantic accuracy of the answer and a penalty for computational cost, specifically the number of LLM calls made. This design is inspired by cost-aware RL frameworks Peters and Schaal (2008).
2.2 Dynamic Tree Construction via RL Agent
@@
2.3 Deep Q-Learning Architecture
We employ Deep Q-Networks (DQN) Mnih et al. (2013) to approximate the optimal policy $\pi^{*}$ that selects actions maximizing the expected reward. The Q-value function $Q(s,a)$ is learned using experience replay and a target network for stable updates.
@@ 2.4 Model Variants
2.4.1 Greedy Hand-Crafted Baseline
@@
2.4.2 DQN with Question-Only State
@@
2.4.3 Transformer-based DQN
Instead of an MLP, this variant uses a Transformer encoder to process the same state vector used in 2.4.2, allowing for potential future extensions to sequence-based state tracking.
2.4.4 DQN with Question + CB + OB
@@
2.4.5 DQN with Reformulation Actions
@@
2.4.6 DQN with Resampling Action
@@
For a more comprehensive comparison of state features across these model variants, please see Appendix C, Table 6.
2.5 Training Regimes
@@
Additionally, an exploration-focused regime and several configurations for reformulation and resampling were tested.
2.6 Summary of Architectures
Table 1 summarizes the different reinforcement learning variants discussed in this paper. The table highlights the models, their state complexity, action space, and any additional relevant notes.
@@ 2.6 Summary of Architectures
Table 1: Summary of reinforcement learning variants.
3 Experimental Setup
@@ 3.1 Datasets
  • HotpotQA Yang et al. (2018) – Multi-hop question answering with sentence-level supporting facts.
  • Musique Trivedi et al. (2022) – Complex multi-hop QA dataset with multiple reasoning steps.
  • 2WikiMultiHopQA (2Wiki) Ho et al. (2020) – Diverse multi-hop QA over Wikipedia.
3.2 Evaluation Metrics
@@
3.3 Training and Testing Configurations
@@
4 Evaluation and Results
4.1 LLM Setup
@@
4.2 In-domain Results
4.2.1 In-domain Accuracy Comparison Across Datasets
We consolidate the in-domain evaluation results across the three datasets—HotpotQA, 2Wiki, and Musique—by including each model’s performance under all three reward configurations. Table 2 presents a comprehensive view of how different models adapt across datasets and reward priorities.
@@ 4.2.1 In-domain Accuracy Comparison Across Datasets
Table 2: In-domain accuracy (%) across reward configurations and datasets. Columns labeled "Orig." represent the results from models in their "Original Configurations" (e.g., Original ProbTree, Greedy Solver, Random Agent), which do not use the reward setups (High, Balanced, Efficient). Remaining columns represent RL-based models under the reward configurations. The best accuracy for each dataset (across Original Config and RL-based Configs) is bolded. - indicates configurations that are not applicable.
The results in Table 2 provide a unified comparison of in-domain performance across models and datasets. While static baselines like ProbTree and the Greedy solver remain strong on HotpotQA and 2Wiki, they underperform significantly on Musique, suggesting that dataset structure plays a critical role in reasoning difficulty.
@@
4.2.2 Accuracy-Cost Tradeoff Analysis
Balancing accuracy and computational cost is crucial for real-world deployment, where LLM calls dominate the cost. Table 2 shows individual accuracy metrics, but Figure 3 visualizes the tradeoff across datasets. Each point represents a (model, configuration, dataset) tuple, with LLM calls on the x-axis and accuracy on the y-axis. Models are differentiated by color, configurations by marker shapes, and datasets by subplot groups.
@@ 4.2.2 Accuracy-Cost Tradeoff Analysis
  • Static Baselines: ProbTree consistently incurs the highest cost (900 calls) while achieving competitive accuracy on 2Wiki (74.6%) and HotpotQA (71.4%) but underperforming on Musique. In contrast, Greedy Solver is more cost-efficient but exhibits dataset-dependent performance, excelling on 2Wiki (77.0%) but lagging on others.
@@
4.3 Results: Cross-Domain Generalization
To evaluate the generalizability of our RL-based approaches, we conducted experiments where models trained on one dataset were tested on other unseen datasets, analyzing their performance and adaptability. Table 2 summarizes the in-domain accuracy, serving as a benchmark, while Table 3 highlights the best accuracy observed for each cross-domain train-test pair.
@@ 4.3 Results: Cross-Domain Generalization
  • Greedy Dominates Across All Cases: The Greedy configuration achieves the best accuracy across all train-test scenarios, establishing itself as the overall top-performing approach. For example, Greedy achieves accuracies of 69.20% for Musique → HotpotQA and 73.00% for Musique → 2Wiki. This significant outperformance sets a strong upper bound for comparison and benchmarks.
@@
Table 3: Best generalization results across train-test pairs. For each pair, we report: (1) the best in-domain accuracy achieved on the test dataset, (2) the best cross-domain result when training on the specified train dataset, and (3) the corresponding model and reward configuration used in the cross-domain setting.
@@ 36.60 Resampling Balanced
4.4 Results: Multi-Domain Generalization
Table 4 summarizes the results for models trained on all three datasets and tested individually on HotpotQA, 2Wiki, and Musique. The performance of the models is evaluated across various RL-based approaches and configurations, as well as baselines such as the Original ProbTree and Greedy Solver.
Table 4: Multi-Domain Generalization Results: Accuracy (%) across Datasets and Approaches. Bold values indicate the best performance within each dataset.
@@ 4.4 Results: Multi-Domain Generalization
On 2Wiki, resampling achieves the highest accuracy at 77.60% (Balanced configuration), marginally surpassing the in-domain best result of 75.60%, as shown in Table 2. This highlights the ability of resampling to generalize effectively under the multi-domain setup.
@@
4.5 Method Usage Analysis
@@
Additional breakdowns by model variant are provided in Appendix B.
5 Related Work
5.1 Structured Prompting and Reasoning Topologies
Recent advancements in structured prompting have introduced various reasoning topologies to enhance the capabilities of large language models (LLMs). Besta et al. Besta et al. (2025) provide a comprehensive taxonomy of these strategies, categorizing them along three dimensions: Topology Scheme, Reasoning Schedule, and AI Pipeline involvement.
Topology Scheme: This dimension classifies the structural format of reasoning, such as chains, trees, or graphs. Our approach aligns with the Tree-of-Thoughts Yao et al. (2023) paradigm, where reasoning is structured hierarchically. However, unlike static trees, our method employs a dynamic tree construction guided by a reinforcement learning (RL) agent, allowing for adaptive reasoning paths based on the problem context.
@@ 5.1 Structured Prompting and Reasoning Topologies
5.2 Adaptive Reasoning Frameworks
Building upon structured prompting, adaptive reasoning frameworks have been proposed to allow LLMs to adjust their reasoning strategies dynamically. Pandey et al. Pandey et al. (2025) introduce the Adaptive Graph of Thoughts (AGoT) framework, which constructs a dynamic directed acyclic graph (DAG) of interdependent reasoning steps at test time. AGoT recursively decomposes complex queries into structured subproblems, selectively expanding only those that require further analysis, thereby unifying the strengths of chain, tree, and graph paradigms into a cohesive framework that allocates computation where it is most needed.
@@ 5.2 Adaptive Reasoning Frameworks
5.3 Self-Guided Multimodal Reasoning
In the realm of multimodal reasoning, self-guided strategies have been explored to enhance LLM performance. Hu et al. Hu et al. (2025) propose the Socratic Questioning framework, where the model engages in self-questioning to guide its reasoning process, particularly in multimodal contexts. This approach heuristically guides multimodal LLMs to focus on relevant visual clues, reducing hallucinations and enhancing the model’s ability to describe fine-grained image details.
@@ Appendix A Complete Cross-Domain Generalization Results
Table 5 presents the full results across all configurations and feature sets for cross-domain generalization experiments. Since the Greedy configuration achieves the best performance across all cases, we highlight it in bold. Additionally, for comparison, we also bold the second-best result among the RL-based configurations (excluding Greedy) for each train-test pair to indicate the next best-performing model in terms of accuracy.
Table 5: Detailed cross-domain generalization results: Accuracy (%) across train-test pairs. Rows include all feature sets and configurations.
@@ Appendix B Method Usage by Model Variant
  • Question-Only agents rely heavily on CB and OB, with occasional Child selections, suggesting that in the absence of additional information (like prior method confidence), the agent defaults to simpler strategies.
  • Question + CB + OB models show a similar pattern but exhibit slightly more confident exploration through Child, likely due to improved state representations derived from answer log-probabilities.
  • Reformulation-based agents display a more balanced distribution across original and reformulated variants, indicating that the additional action space allows for more nuanced strategies. CB_REFORMULATE and OB_REFORMULATE are used significantly more than in other models.
  • Transformer-based models show a strong preference for Child, especially in HotpotQA, aligning with their superior accuracy in that domain. This suggests these models learn richer representations and can better handle deep decomposition.
  • Resampling models are highly skewed toward the ResampleTreeDecompositionThenSolve action, as expected. Their high usage across datasets confirms their behavior of revisiting and revising poor decompositions to enhance performance.

### v3

# From Roots to Rewards: Dynamic Tree Reasoning with Reinforcement Learning

@@ Abstract
Modern language models address complex questions through chain-of-thought (CoT) reasoning and retrieval augmentation, yet struggle with error propagation and knowledge integration. Tree-structured reasoning methods, particularly the Probabilistic Tree-of-Thought (ProbTree) framework, mitigate these issues by decomposing questions into hierarchical structures and selecting answers through confidence-weighted aggregation of parametric and retrieved knowledge. However, ProbTree’s static implementation introduces two key limitations: (1) the reasoning tree is fixed during the initial construction phase, preventing dynamic adaptation to intermediate results, and (2) each node requires exhaustive evaluation of all possible solution strategies, creating computational inefficiency.
We present a dynamic reinforcement learning framework that transforms tree-based reasoning into an adaptive process. Our approach incrementally constructs the reasoning tree based on real-time confidence estimates, while learning optimal policies for action selection (decomposition, retrieval, or aggregation). This maintains ProbTree’s probabilistic rigor while improving both solution quality and computational efficiency through selective expansion and focused resource allocation.
@@
Index Terms:
Reinforcement learning, Tree-of-Thought, Question answering, Large language models, Probabilistic reasoning.
I Introduction
Large language models (LLMs) augmented with tree-structured reasoning, such as ProbTree [1], have shown promise in answering knowledge-intensive complex questions by decomposing queries into sub-questions and probabilistically combining parametric and retrieved knowledge [2].
The ProbTree framework tackles complex question answering by constructing a tree-structured reasoning process. At the root of the tree, the input question is decomposed into smaller sub-questions, forming hierarchical nodes in the reasoning tree. At each node, the framework employs three possible solution strategies: (1) Closed-Book (CB), which leverages parametric knowledge from the language model alone; (2) Open-Book (OB), relying on external retrieval [3] to augment the language model’s outputs; and (3) Child-aggregation (CA), where answers are recursively constructed by combining the outputs of child nodes in the tree. Each node computes a probabilistic confidence score for the outputs of these strategies, and the highest-confidence answer is selected. This selection process ensures that responses are grounded in both parametric (language model) and non-parametric (retrieved) knowledge, making ProbTree well-suited for handling knowledge-intensive queries. However, this static pipeline—where the full reasoning tree and strategy choices are decided in a single step—limits flexibility and introduces computational inefficiencies. Moreover, its brute-force strategy of evaluating all possible solutions per node (Closed-Book, Open-Book, and Child-aggregation) introduces computational inefficiency, especially for deep or wide trees. [^1] Code available at: https://github.com/ahmedehabb/From-Roots-to-Rewards-Dynamic-Tree-Reasoning-with-RL
I-A Motivation
@@ I-C Contributions
  • Empirical validation showing competitive performance with reduced computational cost on standard benchmarks (HotpotQA [4], MusiQue [5], 2WikiMultihopQA [6])
@@
II Reinforcement Learning Framework for Dynamic Tree-of-Thought Reasoning
To overcome the limitations of static and brute-force tree generation in the original ProbTree framework [1], we introduce a reinforcement learning (RL) based agent that dynamically generates reasoning trees and selects actions at each step to answer knowledge-intensive questions efficiently and accurately.
II-A Problem Formulation
We formulate the decision process of tree construction and answer selection as a Markov Decision Process (MDP) [7], where:
  • State ($s$): Encodes information about the current node in the tree, including semantic features, structural position, confidence scores, and answer embeddings [8].
@@ II-A Problem Formulation
  • Reward ($r$): A combination of semantic accuracy of the answer and a penalty for computational cost, specifically the number of LLM calls made. This design is inspired by cost-aware RL frameworks [9].
II-B Dynamic Tree Construction via RL Agent
@@
II-C Deep Q-Learning Architecture
We employ Deep Q-Networks (DQN) [10] to approximate the optimal policy $\pi^{*}$ that selects actions maximizing the expected reward. The Q-value function $Q(s,a)$ is learned using experience replay and a target network for stable updates.
@@ II-D Model Variants
II-D1 Greedy Hand-Crafted Baseline
@@
II-D2 DQN with Question-Only State
@@
II-D3 Transformer-based DQN
Instead of an MLP [11], this variant uses a Transformer encoder to process the same state vector used in II-D2, allowing for potential future extensions to sequence-based state tracking.
II-D4 DQN with Question + CB + OB
@@
II-D5 DQN with Reformulation Actions
@@
II-D6 DQN with Resampling Action
@@
II-E Training Regimes
@@
II-F Summary of Architectures
Table I summarizes the different reinforcement learning variants discussed in this paper. The table highlights the models, their state complexity, action space, and any additional relevant notes.
TABLE I: Summary of reinforcement learning variants.
@@ II-F Summary of Architectures
III Experimental Setup
@@ III-A Datasets
  • HotpotQA [4] – Multi-hop question answering with sentence-level supporting facts.
  • Musique [5] – Complex multi-hop QA dataset with multiple reasoning steps.
  • 2WikiMultiHopQA (2Wiki) [6] – Diverse multi-hop QA over Wikipedia.
III-B Evaluation Metrics
@@
III-C Training and Testing Configurations
@@
IV Evaluation and Results
IV-A LLM Setup
@@
IV-B In-domain Results
IV-B1 In-domain Accuracy Comparison Across Datasets
We consolidate the in-domain evaluation results across the three datasets—HotpotQA, 2Wiki, and Musique—by including each model’s performance under all three reward configurations. Table II presents a comprehensive view of how different models adapt across datasets and reward priorities.
TABLE II: In-domain accuracy (%) across reward configurations and datasets. Columns labeled ”Orig.” represent the results from models in their ”Original Configurations” (e.g., Original ProbTree, Greedy Solver, Random Agent), which do not use the reward setups (High, Balanced, Efficient). Remaining columns represent RL-based models under the reward configurations. The best accuracy for each dataset (across Original Config and RL-based Configs) is bolded. - indicates configurations that are not applicable.
@@ IV-B1 In-domain Accuracy Comparison Across Datasets
The results in Table II provide a unified comparison of in-domain performance across models and datasets. While static baselines like ProbTree and the Greedy solver remain strong on HotpotQA and 2Wiki, they underperform significantly on Musique, suggesting that dataset structure plays a critical role in reasoning difficulty.
@@
IV-B2 Accuracy-Cost Tradeoff Analysis
Balancing accuracy and computational cost is crucial for real-world deployment, where LLM calls dominate the cost. Table II shows individual accuracy metrics, but Figure 3 visualizes the tradeoff across datasets. Each point represents a (model, configuration, dataset) tuple, with LLM calls on the x-axis and accuracy on the y-axis. Models are differentiated by color, configurations by marker shapes, and datasets by subplot groups.
@@ IV-B2 Accuracy-Cost Tradeoff Analysis
  • Static Baselines: ProbTree consistently incurs the highest cost (900 calls) while achieving competitive accuracy on HotpotQA (71.4%) and 2Wiki (74.6%) but underperforming on Musique (44.8%). In contrast, Greedy Solver is substantially more cost-efficient: it outperforms all methods on 2Wiki (77.0%), remains competitive on HotpotQA (70.4%), but lags behind on Musique (44.2%). Random Agent performs the worst across all datasets.
@@
IV-C Results: Cross-Domain Generalization
To evaluate the generalizability of our RL-based approaches, we conducted experiments where models trained on one dataset were tested on other unseen datasets, analyzing their performance and adaptability. Table II summarizes the in-domain accuracy, serving as a benchmark, while Table III highlights the best accuracy observed for each cross-domain train-test pair.
@@ IV-C Results: Cross-Domain Generalization
  • Greedy Provides the Strongest Baseline: The Greedy configuration consistently achieves the highest or near-highest cross-domain accuracy across train–test scenarios, establishing itself as the most reliable static baseline. For example, it reaches 69.20% for Musique→HotpotQA and 73.00% for Musique→2Wiki.
@@
TABLE III: Best generalization results across train–test pairs. For each test dataset, we report: (1) its best in-domain accuracy (upper bound), and for each train–test pair we provide two cross-domain results: (a) the Greedy Solver in its original configuration, serving as the strongest static baseline, and (b) the best-performing RL-based model under any reward configuration.
@@ 36.60 Resampling Balanced
IV-D Results: Multi-Domain Generalization
Table IV summarizes the results for models trained on all three datasets and tested individually on HotpotQA, 2Wiki, and Musique. The performance of the models is evaluated across various RL-based approaches and configurations, as well as baselines such as the Original ProbTree and Greedy Solver.
TABLE IV: Multi-Domain Generalization Results: Accuracy (%) across Datasets and Approaches. Bold values indicate the best performance within each dataset.
@@ IV-D Results: Multi-Domain Generalization
On 2Wiki, resampling achieves the highest accuracy at 77.60% (Balanced configuration), marginally surpassing the in-domain best result of 75.60%, as shown in Table II. This highlights the ability of resampling to generalize effectively under the multi-domain setup.
@@
IV-E Method Usage Analysis
@@
V Related Work
V-A Structured Prompting and Reasoning Topologies
Recent advancements in structured prompting have introduced various reasoning topologies to enhance the capabilities of large language models (LLMs). Besta et al. [12] provide a comprehensive taxonomy of these strategies, categorizing them along three dimensions: Topology Scheme, Reasoning Schedule, and AI Pipeline involvement.
Topology Scheme: This dimension classifies the structural format of reasoning, such as chains [13], trees [14], or graphs. Our approach aligns with the Tree-of-Thoughts [14] paradigm, where reasoning is structured hierarchically. However, unlike static trees, our method employs a dynamic tree construction guided by a reinforcement learning (RL) agent, allowing for adaptive reasoning paths based on the problem context.
@@ V-A Structured Prompting and Reasoning Topologies
V-B Adaptive Reasoning Frameworks
Building upon structured prompting, adaptive reasoning frameworks have been proposed to allow LLMs to adjust their reasoning strategies dynamically. Pandey et al. [15] introduce the Adaptive Graph of Thoughts (AGoT) framework, which constructs a dynamic directed acyclic graph (DAG) of interdependent reasoning steps at test time. AGoT recursively decomposes complex queries into structured subproblems, selectively expanding only those that require further analysis, thereby unifying the strengths of chain, tree, and graph paradigms into a cohesive framework that allocates computation where it is most needed.
@@ V-B Adaptive Reasoning Frameworks
V-C Self-Guided Multimodal Reasoning
In the realm of multimodal reasoning, self-guided strategies have been explored to enhance LLM performance. Hu et al. [16] propose the Socratic Questioning framework, where the model engages in self-questioning to guide its reasoning process, particularly in multimodal contexts. This approach heuristically guides multimodal LLMs to focus on relevant visual clues, reducing hallucinations and enhancing the model’s ability to describe fine-grained image details.
@@ Appendix A Complete Cross-Domain Generalization Results
Table V presents the full results across all configurations and feature sets for cross-domain generalization experiments. Since the Greedy configuration achieves the best performance across all cases, we highlight it in bold. Additionally, for comparison, we also bold the second-best result among the RL-based configurations (excluding Greedy) for each train-test pair to indicate the next best-performing model in terms of accuracy.
TABLE V: Detailed cross-domain generalization results: Accuracy (%) across train-test pairs. Rows include all feature sets and configurations.
@@ Appendix B Method Usage by Model Variant
  • Question-Only agents rely heavily on CB and OB, with occasional Child selections, defaulting to simpler strategies.
  • Question + CB + OB models show a similar pattern but with slightly more Child usage from improved state representations derived from answer log-probabilities.
  • Reformulation-based agents display a more balanced distribution across original and reformulated variants, with CB_REFORMULATE and OB_REFORMULATE used significantly more, indicating that the additional action space allows for more nuanced strategies.
  • Transformer-based models strongly prefer Child, especially in HotpotQA, aligning with their superior accuracy in that domain.
  • Resampling models are skewed toward the ResampleTreeDecompositionThenSolve action, confirming their tendency to revise poor decompositions to enhance performance.